OpenStudy (anonymous):

The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 90 feet with an initial velocity of 50 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. Part B: What is the maximum height that the projectile will reach? Show your work. Part C: Another object moves in the air along the path of g(t) = 28 + 48.8t where g(t) is th

OpenStudy (anonymous):

A. H(t)=-16t^2+v t+s s=90 v=50 ft/sec when the projectile is on the ground H=0 \[0=-16 t^2+50 t+90,find ~t\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

B. again \[H=-16t^2+50t+90,\frac{ dH }{ dt }=-32 t+50\] \[when~ the~ body ~is ~at ~maximum~ height \frac{ dH }{ dt }=0\] find t and then H at this t.

OpenStudy (anonymous):

this is to confusing for me :/

OpenStudy (anonymous):

for part B initial velocity v=50 given when the projectile is at maximum height its velocity becomes zero ,then it comes back to ground.

OpenStudy (anonymous):

\[\frac{ dH }{ dt }\] gives the velocity of projectile at any time t.

OpenStudy (anonymous):

-32t+50=0 \[t=\frac{ 50 }{ 32 }=\frac{ 25 }{ 16 } \sec\] find H at t=25/16 sec